OpenStudy (anonymous):

What is the slope of the line that passes through (1, 4) and (1, -3)? give a medal to first that answers

OpenStudy (anonymous):

i think it is undefined

OpenStudy (anonymous):

kjos what are you typing

OpenStudy (whpalmer4):

Yes, the slope is undefined, because there are multiple points for the same x value. If you used the formula for the slope given two points, you would be dividing by 0.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

we can describe a line by the equation \[ ax + by = c \] . For a line that passes through point (x1, y1), the equation would be \[ ax_1 + by_1 = c \] and for the line that passes through (x2, y2), the equation would be \[ax_2 + by_2 = c \] Now, if we subtract equation 3 from 2 we will get \[ a(x_1-x_2) + b(y_1-y_2) = 0\] Thus, the equation of the line that passes through those two points is: a(0) + 7b =0, and thus b = 0, and thus, the equation is ax=c i.e. x = c/a i.e. it is a vertical line and thus the slope is infinity